OpenStudy (anonymous):

If f(x) = 3x 2 - x, find f(-2). 10 14 38

OpenStudy (anonymous1):

\[f(x)=3x^2-x\] To find f(-2) just replace x with -2: \[f(-2)=3(-2)^2-(-2)\] \[f(-2)=3*4+2=12+2\] \[f(-2)=14\]